OpenStudy (anonymous):

Find the exact value by using a half-angle identity. cos{5pi/12}

OpenStudy (dumbcow):

half-angle identity \[\cos(\theta/2) = \pm \sqrt{\frac{1+\cos \theta}{2}}\] theta = 5pi/6

OpenStudy (dumbcow):

find cos(5pi/6)

OpenStudy (anonymous):

-\[-\sqrt{3}/2\]

OpenStudy (dumbcow):

yep, now plug that in and simplify